# Yong Ji1, Junye Li1 and Rui Yang∗2,3,4

### v1

@@
Abstract. Saturated sets and its reduced case the set of generic points are two types of important fractal-like sets in Multifractal analysis of dynamical systems. In the context of infinite entropy systems, this paper aims to give some qualitative aspects of saturated sets and the set generic points in both topological and measure-theoretic situations. For systems with specification property, we establish the certain variational principles for saturated sets in terms of Bowen and packing metric mean dimensions, and show the upper capacity metric mean dimension of saturated sets have full metric mean dimension. All results are useful for understanding the topological structures of dynamical systems with infinite topological entropy.
@@
Keywords and phrases: Metric mean dimension; Saturated sets; Generic point; Variational principle; Multifractal analysis
@@ 1 Introduction
Let $(X,d,f)$ be a topological dynamical system (t.d.s. for short), where $(X,d)$ is a compact metric space with metric $d$, and $f:X\to X$ is a homeomorphism. Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(x)$ denote the sets of Borel probability measures endowed with weak-topology, $f$-invariant Borel probability measures, and ergodic $f$-invariant Borel probability measures on $X$, respectively. Measure-theoretic entropy and topological entropy are two vital invariants for us to understand the topological complexity of dynamical systems, which are related by the classical variational principle:
@@
The multifractal analysis, which can date back to Besicovitch, considers the decomposition of the whole phase space into smaller subset consisting of the points with the similar dynamical behaviors, or known as multifractal spectrum, and then from the geometrical or topological viewpoint describes the “size” of those subsets by using fractal dimensions and dimensional entropies. Such a procedure allows us to get more information about the dynamics and even recover the partial dynamical behaviors of the system, which we call such phenomenon multifractal rigidity. We aim to study the multifractal analysis of saturated sets in infinite entropy systems.
Let us state some backgrounds about the dimensional entropies of saturated sets. Given $x\in X$, by $\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}\in M(X)$, where $\delta_{x}$ is the Dirac mass at $x$, we denote the $n$-th empirical measure of the Birkhoff average w.r.t. $x$. Then the limit-point set $V_{f}(x)$ of $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a nonempty compact connected subset of $M_{f}(X)$ in the sense of weak-topology (cf. [DGS76, Proposition 3.8]). Specially, when $V_{f}(x)=\{\mu\}$ for some measure $\mu$, one says that $x$ is a generic point of $\mu$. The set of all generic points of $\mu$ is denoted by $G_{\mu}$. Then the Birkhoff ergodic theorem implies that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic. Bowen [Bow73] showed for any ergodic measure $\mu$, the Bowen topological entropy of $G_{\mu}$ is equal to $h_{\mu}(f)$, that is,
@@
For non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$ and even easily constructs the example showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$ such that Bowen’s result is not valid. It is natural to expect whether there are some proper conditions such that the analogous equality still holds for non-ergodic measures. A subset $D$ of $X$ is called a saturated set if $x\in D$ and $V_{f}(x)=V_{f}(y)$, then $y\in D$. More generally, given a nonempty closed connected subset $K\subset M_{f}(X)$, this defines a saturated set $G_{K}=\{x\in X:V_{f}(x)=K\}$ associated with $K$. For systems with specification property, including hyperbolic systems, Sigmund [Sig72, Sig74] firstly proved the existence of the saturated sets, which was later extended to non-uniformly hyperbolic and non-uniformly expanding systems [LST13, TV17]. As for the estimation of Bowen topological entropy of saturated sets, for systems with $g$-almost product property that is weaker than specification property, which solely requires a point partially $\epsilon$-shadowing the given finite pieces of orbit-segments, and uniform separation property (e.g. expansive and more generally asymptotically $h$-expansive systems), Pfister and Sullivan [PS07] established the certain variational principle for any compact connected non-empty set $K\subset M_{f}(X)$,
@@
holds for any $\mu\in M_{f}(X)$, which extends Bowen’s remarkable result to any non-ergodic measures. Consequently, the work of [Bow73, PS07] leads to an intensive study toward the estimation of different types of dimensional entropies of saturated sets. Packing entropy of non-compact subset was introduced by Feng and Huang [FH12] in a way resembling the definition of packing dimension. Under a stronger condition of specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] further established the following variational principle for any compact connected non-empty set $K\subset M_{f}(X)$,
@@
where $h_{top}^{P}(f,K)$ is the packing entropy of $K$. Later, weakening the specification property, for systems with $g$-almost product property, Hou, Tian and Zhang [HTZ23] not only proved that the variational principle $h_{top}^{P}(f,G_{K}^{C})=\sup\{h_{\mu}(f):\mu\in K\}$ holds for (mixed) saturated sets with more complicated dynamics, but also showed the upper capacity topological entropy of the saturated sets have full topological entropy, i.e.
@@
where $G_{K}^{C}:=G_{K}\cap\{x\in X:C_{f}(X)\subset\omega_{f}(x)\},$ $\omega_{f}(x):=\bigcap_{n=0}^{\infty}\overline{\bigcup_{k=n}^{\infty}\{f^{k}x\}}$ is the $\omega$-limit set of $x$, and $C_{f}(X)$ is the measure center given by $C_{f}(X):=\overline{\cup_{\mu\in M_{f}(X)}{\rm supp}{(\mu)}}$, ${\rm supp}(\mu)$ is the supported set of $\mu$. Besides, for systems satisfying $g$-almost product property and uniform separation property, Huang, Tian and Wang [HTW19] showed Pfister and Sullivan’s variational principle still hold for transitive-saturated set $G_{K}^{T}:=G_{K}\cap{\rm Tran}$, where ${\rm Tran}$ denotes the set of transitive points. Readers can turn to [JCZ22, LLS17] for more results involving the dimensional entropies of saturated sets in nonuniformly hyperbolic systems, and [ZC18, DZZ23, Z18, RTZ23, ZRZ23] for some results about the amenable Bowen and packing topological entropies of saturated sets in the context of the actions of amenable groups.
According to the possible values of topological entropy, dynamical systems can be divided into three types: zero entropy systems, finite positive entropy systems and infinite entropy systems. Infinite entropy system always means the system may possess quiet complicated topological complexity, so the notion of topological entropy no longer provides more information about the dynamics of the system. In 2000, Lindenstrauss and Weiss [LW00] introduced a refined notion for topological entropy, called metric mean dimension, which is a dynamical analogue of box dimension to measure how fast the $\epsilon$-topological entropy diverges to $\infty$ as $\epsilon$ goes to 00. It turns out that any finite entropy system has zero metric mean dimension. See some applications of metric mean dimension in the analog compression of information theory [GS20], and in estimating the upper bound of mean dimension of some complex systems [Tsu18a, Tsu18b]. The notion of measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, dues to Gutman and Śpiewak[GS21]. More precisely, the lower and upper mean Rényi information dimensions of $\mu\in M_{f}(X)$ are respectively defined by
@@
which can be explained how fast the Kolmogorov-Sinai $\epsilon$-entropy $\inf_{\mathrm{diam}\xi<\epsilon}\limits h_{\mu}(f,\xi)$ diverges to $\infty$ as $\epsilon\to 0$. Some variational principles of metric mean dimension of $\mathbb{Z}$-actions can be found in [LW00, GS21, S22, YCZ23a]. Then the aforementioned work suggests the following questions for infinite entropy systems:
@@
For (1), the authors in [YCZ24] showed the packing metric mean dimension of the set of generic points equals to the mean Rényi information dimension, however it holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula in infinite entropy systems (cf.[YCZ24, §5, Question,(1)]); for (2), noting the uniform separation property implies the system has finite topological entropy, so extending the work of [PS07] to the setting of infinite entropy systems encounters some nontrivial challenges. Only the assumption of $g$-almost product property, Yuan [Y23] investigated the Bowen metric mean dimension of $G_{K}$, and showed that for any $\{\mu_{1},\mu_{2},...,\mu_{k}\}\subset M_{f}(X)$ and any non-empty compact connected set $K$ containing in the convex closure ${\rm cov}\{\mu_{1},...,\mu_{k}\}$,
@@
However, for systems with specification property we can give a full answer to question (2). Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the following statements of main results involve more general (mixed) saturated sets. For simplicity, we write
@@
Different types of metric mean dimensions of saturated sets exhibit different dynamical behaviors. There exist certain variational principles for Bowen and packing metric metric mean dimensions of saturated sets (Theorems 1.1 and 1.2), while the upper capacity metric mean dimension of saturated sets has full metric mean dimension (Theorem 1.4).
@@
Inspired the work of [GS21, CPV24], we define two variants of mean Rényi information dimension of invariant measures. For any closed set $K$ of $M_{f}(X)$ and $\mu\in K$, by $M_{K}(\mu)$ we denote the space of sequences of Borel probability measures in $K$ which converge to $\mu$ in the weak-topology. We respectively define the lower and upper mean Rényi information dimensions of $\mu\in K$ as
@@
We remark that Theorem 1.3 does not need the connectedness of $K$ since any convex set in a topological vector space is automatically connected.
@@
As a direct consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case is $K$ being just a single point. The Corollary 1.5 presents the metric mean dimensions of generic points of any invariant measures for systems with specification property, which extends the result given in [YCZ24, Theorem 1.1] to all non-ergodic measures.
@@
The “Consequently” part can realize for some dynamical systems. For example, the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$ with the compatible product metric $d(x,y)=\sum_{n\in\mathbb{Z}}\frac{||x_{n}-y_{n}||_{\mathbb{R}^{m}}}{2^{|n|}}$, and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is product measure of the Lebesgue measure $\mathcal{L}$ on $[0,1]^{m}$. The concept of maximal metric mean dimension measure was proposed in [YCZ23a], yet for what kind of infinite entropy systems can have maximal metric mean dimension measure is still a difficult problem. See [LR24] for the existence of maximal metric mean dimension measure for typical conservative homeomorphisms. For some certain systems, by Gutman- Śpiewak variational principle (see Theorem 2.4) and Corollary 1.5 we have following (abstract) criterion for maximal metric mean dimension measure for systems with specification property.
@@
    There are plenty of infinite entropy systems with specification property, for instance, the usual the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$, and the examples presented in [BR23, Section IV] such that our main results can be applied to such systems.
@@
    We believe it is trivial to extend the aforementioned results to metric mean dimensions with potential by considering some modifications of the present proofs. The novelty of our work not only presents a proper and full extension of the previous work in [PS07, ZCC12, HTW19] to the setting of infinite entropy systems of $\mathbb{Z}$-actions with specification property, but also may be helpful for the multifractal analysis of other types of fractal-like sets in infinite entropy systems. It remains unclear whether there are some proper substitutions of the uniform separation property so that all results are true for $g$-almost product property, or beyond $\mathbb{Z}$-actions.
The paper is organized as follows. In Section 2, we recall the precise definitions of metric mean dimensions and the associated properties, and two types of measure-theoretic $\epsilon$-entropies, and the specification property. In section 3, we prove the main results. In Section 4, we give two applications of main results. In section 5, we show the metric mean dimensions of generic points of full shift over compact metric spaces.
@@ 2 Preliminaries
In this section, we recall the concepts of metric mean dimensions in both topological and measure-theoretic situations, and the precise definition of specification property in dynamical systems.
@@ 2.1 Three types of metric mean dimensions
Let $(X,d,f)$ be a dynamical system. For $n\in N$, we define the Bowen metric $d_{n}$ on $X$ as
@@
$h_{top}(f,Z,\epsilon)=\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(Z,\epsilon).$
Analogous to the definition of box dimension in fractal geometry, we can formulate the definition of upper capacity metric mean dimension of subsets using separated sets.
@@
When $Z=X$, we sometimes write $h_{top}(f,\epsilon)=h_{top}^{UC}(f,X,\epsilon)$, $\overline{\rm mdim}_{M}^{UC}(f,X,d)=\overline{\rm mdim}_{M}(f,X,d)$ respectively, then in this case it is exactly reduced to Lindenstrauss and Weiss’s classical notion introduced in [LW00]. Such a formulation allows us to extend the metric mean dimensions to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
Recall that the topological entropy of $f$ is limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 00, that is,
$h_{top}(f)=\lim_{\epsilon\to 0}h_{top}(f,\epsilon)=\sup_{\epsilon>0}h_{top}(f,\epsilon).$
For any system admitting the infinite topological entropy, topological entropy does not provide the additional information of the dynamics except we have known the system has pretty complicated topological complexity. Thus, (upper capacity) metric mean dimension can be viewed as a quantity to detect how fact the $\epsilon$-upper capacity topological entropy diverges to the infinite entropy of the system as $\epsilon$ tends to 00. More precisely, for sufficiently $\epsilon>0$, we may think of
@@
It is easy to see that any finite entropy system has zero metric mean dimension. This further verifies that metric mean dimension is a good candidate to capture the dynamics of infinite systems.
@@
The quantity $M(Z,s,N,\epsilon)$ does not decrease as $N$ increase. So we set
@@
Since the quantity $P(Z,s,N,\epsilon)$ does not increase as $N$ increases, the following limit exists:
@@
One can replace $\limsup_{\epsilon\to 0}$ by $\liminf_{\epsilon\to 0}$ in above definitions to defined the corresponding three types of lower metric mean dimensions on subsets. We summarize some fundamental properties of metric mean dimensions derived in [YCZ22, Proposition 3.4] and [YCZ23b, Proposition 2.6] as follows:
@@
and $h_{top}(f,Z,\epsilon)=\max_{1\leq i\leq N}h_{top}(f,Z_{i},\epsilon)$ if $Z=\bigcup_{i=1}^{N}Z_{i}$.
@@
$(3)$ If $Z$ is an any non-empty subset of $X$, then for every $10<\epsilon<$ 1, we have
$h_{top}^{B}(f,Z,3\epsilon)\leq h_{top}^{P}(f,Z,\epsilon)\leq h_{top}(f,Z,\epsilon).$
@@
Observe that the metric mean dimension of a given subset is in general hard to calculate its precise value. However, the Entropy Distribution Principle allows us to get the lower bounds of the $\epsilon$-Bowen and $\epsilon$-packing topological entropies of some fractal-like sets in multifractal analysis of dynamical systems. Consequently, the powerful tool enables us to get the lower bounds of entropy and metric mean dimension of the fractal-like sets.
@@
[BR23, Lemma 13] Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Let $Z$ be a Borel set of $X$ and $s\geq 0$. If there exists a Borel probability measure $\mu$ and a constant $C>0$ satisfying:
@@
Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Let $Z$ be a Borel set of $X$ and $s\geq 0$. If there exists a Borel probability measure $\mu$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ of positive integer that converges to $\infty$ satisfying:
@@
holds for any Borel subset $E\subset Z$ with $\mu(E)>0$. For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of $E$. Then by $5r$-covering lemma stated in [Mat95, Theorem 2.1], there is a countable pairwise disjoint sub-collection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, and in particular the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint, such that

### v2

@@
Abstract. Saturated set and its reduced case, the set of generic points, constitute two significant types of fractal-like sets in multifractal analysis of dynamical systems. In the context of infinite entropy systems, this paper aims to give some qualitative aspects of saturated sets and the set of generic points in both topological and measure-theoretic perspectives. For systems with specification property, we establish the certain variational principles for saturated sets in terms of Bowen and packing metric mean dimensions, and show the upper capacity metric mean dimension of saturated sets have full metric mean dimension. All results are useful for understanding the topological structures of dynamical systems with infinite topological entropy.
@@
Keywords and phrases: Metric mean dimension; Saturated set; Generic point; Variational principle; Multifractal analysis
@@ 1 Introduction
Let $(X,d,f)$ be a topological dynamical system (t.d.s. for short), where $(X,d)$ is a compact metric space with a metric $d$, and $f:X\to X$ is a homeomorphism. Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(x)$ denote the sets of Borel probability measures endowed with weak-topology, $f$-invariant Borel probability measures, and ergodic $f$-invariant Borel probability measures on $X$, respectively. Measure-theoretic entropy and topological entropy are two vital invariants for us to understand the topological complexity of dynamical systems, which are related by the classical variational principle:
@@
The multifractal analysis, which can date back to Besicovitch, takes into account the decomposition of the whole phase space into smaller subsets consisting of the points with the similar dynamical behaviors(or known as multifractal spectrum), and then from the geometrical or topological viewpoint describes the “size” of those subsets by using fractal dimensions and dimensional entropies. Such a procedure allows us to get more information about the dynamics and even recovers the partial dynamical behaviors of the system, which is called phenomenon multifractal rigidity. We aim to study the multifractal analysis of saturated sets in infinite entropy systems.
Let us mention some relevant results about the dimensional entropies of saturated sets. Given $x\in X$, by $\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}\in M(X)$, where $\delta_{x}$ is the Dirac mass at $x$, we denote the $n$-th empirical measure of the Birkhoff average w.r.t. $x$. Then the limit-point set $V_{f}(x)$ of $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a nonempty compact connected subset of $M_{f}(X)$ in the sense of weak-topology (cf. [DGS76, Proposition 3.8]). Specially, when $V_{f}(x)=\{\mu\}$ for some measure $\mu$, one says that $x$ is a generic point of $\mu$. The set of all generic points of $\mu$ is denoted by $G_{\mu}$. Then the Birkhoff’s ergodic theorem implies that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic. Bowen [Bow73] proved that for any ergodic measure $\mu$, the Bowen topological entropy of $G_{\mu}$ is equal to $h_{\mu}(f)$, that is,
@@
For non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$ and even easily constructs the example showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$ such that Bowen’s result is not valid. It is natural to expect whether there are some proper conditions such that the equality still holds for non-ergodic measures. A subset $D$ of $X$ is called a saturated set if $x\in D$ and $V_{f}(x)=V_{f}(y)$, then $y\in D$. More generally, given a nonempty closed connected subset $K\subset M_{f}(X)$, we can obtain a saturated set $G_{K}=\{x\in X:V_{f}(x)=K\}$ associated with $K$. For systems with specification property, including hyperbolic systems, Sigmund [Sig72, Sig74] firstly proved the existence of the saturated sets, which was later extended to non-uniformly hyperbolic and non-uniformly expanding systems by several authors [LST13, TV17]. As for the estimation of Bowen topological entropy of saturated sets, for systems with $g$-almost product property that is weaker than specification property, which solely requires a point partially $\epsilon$-shadowing the given finite pieces of orbit-segments, and with uniform separation property (e.g. expansive and more generally asymptotically $h$-expansive systems), Pfister and Sullivan [PS07] established the certain variational principle for any nonempty compact connected set $K\subset M_{f}(X)$,
@@
holds for all $\mu\in M_{f}(X)$, which extends Bowen’s remarkable result to non-ergodic measures. Consequently, the work of [Bow73, PS07] leads to an intensive study toward the estimation of different types of dimensional entropies of saturated sets. Packing entropy of non-compact subsets was introduced by Feng and Huang [FH12] in a way resembling the definition of packing dimension. Additionally, under the (stronger) conditions of specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] showed the following variational principle for any nonempty compact connected set $K\subset M_{f}(X)$,
@@
where $h_{top}^{P}(f,K)$ is the packing entropy of $K$. Afterwards, weakening the specification property, for systems with $g$-almost product property, Hou, Tian and Zhang [HTZ23] not only proved that the variational principle $h_{top}^{P}(f,G_{K}^{C})=\sup\{h_{\mu}(f):\mu\in K\}$ holds for (mixed) saturated sets with more complicated dynamics, but also showed the upper capacity topological entropy of the saturated sets have full topological entropy, i.e.
@@
where $G_{K}^{C}:=G_{K}\cap\{x\in X:C_{f}(X)\subset\omega_{f}(x)\},$ $\omega_{f}(x):=\bigcap_{n=0}^{\infty}\overline{\bigcup_{k=n}^{\infty}\{f^{k}x\}}$ is the $\omega$-limit set of $x$, and $C_{f}(X)$ is the measure center given by $C_{f}(X):=\overline{\cup_{\mu\in M_{f}(X)}{\rm supp}{(\mu)}}$, ${\rm supp}(\mu)$ is the supported set of $\mu$. Furthermore, for systems satisfying $g$-almost product property and uniform separation property, Huang, Tian and Wang [HTW19] showed Pfister and Sullivan’s variational principle is valid for transitive-saturated set $G_{K}^{T}:=G_{K}\cap{\rm Tran}$, where ${\rm Tran}$ denotes the set of transitive points. Readers can turn to [JCZ22, LLS17] for more results involving the dimensional entropies of saturated sets in the setting of nonuniformly hyperbolic systems, and [ZC18, DZZ23, Z18, RTZ23, ZRZ23] for some results concerning the amenable Bowen and packing topological entropies of saturated sets in the context of the actions of amenable groups.
According to the possible values of topological entropy, dynamical systems can be divided into three types: zero entropy systems, finite positive entropy systems and infinite entropy systems. Infinite entropy system always means the system may possess quiet complicated topological complexity, so the notion of topological entropy no longer provides more information about the dynamics of the system. In 2000, Lindenstrauss and Weiss [LW00] introduced a refined notion for topological entropy, called metric mean dimension, which is a dynamical analogue of box dimension to measure how fast the $\epsilon$-topological entropy diverges to $\infty$ as $\epsilon$ goes to 00. It turns out that any finite entropy system has zero metric mean dimension. See some applications of metric mean dimension in the analog compression of information theory [GS20], and in estimating the upper bound of mean dimension of some complex systems [Tsu18a, Tsu18b]. The notion of measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, is due to Gutman and Śpiewak[GS21]. More precisely, the lower and upper mean Rényi information dimensions of $\mu\in M_{f}(X)$ are respectively defined by
@@
which can be explained how fast the Kolmogorov-Sinai $\epsilon$-entropy $\inf_{\mathrm{diam}\xi<\epsilon}\limits h_{\mu}(f,\xi)$ diverges to $\infty$ as $\epsilon\to 0$. Some variational principles of metric mean dimension of $\mathbb{Z}$-actions can be found in [LW00, GS21, S22, YCZ23a]. Then the aforementioned work suggests the following questions for saturated sets in infinite entropy systems:
@@
For (1), the authors in [YCZ24] showed the packing metric mean dimension of the set of generic points equals to the mean Rényi information dimension, however it holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula in infinite entropy systems (cf.[YCZ24, §5, Question,(1)]); for (2), noting the uniform separation property implies the system has finite topological entropy, so extending the work of [PS07] to the framework of infinite entropy systems encounters some nontrivial challenges. Only the assumption of $g$-almost product property, Yuan [Y23] investigated the Bowen metric mean dimension of $G_{K}$, and showed that for any $\{\mu_{1},\mu_{2},...,\mu_{k}\}\subset M_{f}(X)$ and any non-empty compact connected set $K$ contained in the convex closure ${\rm cov}\{\mu_{1},...,\mu_{k}\}$,
@@
Prior to the comparison of different types of specification-like property, we only focus on the systems with specification property and give a full answer to question (2). Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the following statements of main results involve more general (mixed) saturated sets. For simplicity, we write
@@
Saturated sets exhibit different dynamical behaviors in terms of different types of metric mean dimensions. There exist certain variational principles for Bowen and packing metric metric mean dimensions of saturated sets (Theorems 1.1 and 1.2), while the upper capacity metric mean dimension of saturated set has full metric mean dimension (Theorem 1.4).
@@
Inspired by the work of [GS21, CPV24], we define two variants of mean Rényi information dimension of invariant measures. For any closed set $K$ of $M_{f}(X)$ and $\mu\in K$, by $M_{K}(\mu)$ we denote the space of sequences of Borel probability measures in $K$ which converge to $\mu$ in the weak-topology. We respectively define the (modified) lower and upper mean Rényi information dimensions of $\mu\in K$ as
@@
We remark that Theorem 1.3 does not require the connectedness of $K$ since any convex set in a topological vector space is automatically connected.
@@
As a direct consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case is $K$ being exactly a single point. The Corollary 1.5 presents the metric mean dimensions of generic points of any invariant measures for systems with specification property, which extends the result given in [YCZ24, Theorem 1.1] to all non-ergodic measures.
@@
The “Consequently” part can be realized for some dynamical systems. For example, the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$ with the compatible product metric $d(x,y)=\sum_{n\in\mathbb{Z}}\frac{||x_{n}-y_{n}||_{\mathbb{R}^{m}}}{2^{|n|}}$, and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is product measure of the Lebesgue measure $\mathcal{L}$ on $[0,1]^{m}$. The concept of maximal metric mean dimension measure was proposed in [YCZ23a], yet for what kind of infinite entropy systems can have maximal metric mean dimension measure is still a difficult problem. See [LR24] for the existence of maximal metric mean dimension measure for typical conservative homeomorphisms. By Gutman- Śpiewak variational principle (see Theorem 2.4) and Corollary 1.5 we have following (abstract) criterion for maximal metric mean dimension measure for systems with specification property.
@@
    There are plenty of infinite entropy systems with specification property, for instance, the usual full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$, and the examples presented in [BR23, Section IV] such that our main results can be applied to such systems.
@@
    We believe that it is trivial to extend the aforementioned results to metric mean dimensions with potential by considering some modifications of the present proofs. The novelty of our work not only presents a proper and full extension of the previous work in [PS07, ZCC12, HTW19] to the setting of infinite entropy systems of $\mathbb{Z}$-actions with specification property, but also may be helpful for the multifractal analysis of other types of fractal-like sets in infinite entropy systems. It remains unclear whether there are some proper substitutions of the uniform separation property so that all results are true for $g$-almost product property, or beyond $\mathbb{Z}$-actions.
The paper is organized as follows. In section 2, we recall the precise definitions of metric mean dimensions and the associated properties, and two types of measure-theoretic $\epsilon$-entropies, and the specification property. In section 3, we prove the main results. In section 4, we give two applications of the main results. In section 5, we show the metric mean dimensions of generic points of full shift over compact metric spaces.
@@ 2 Preliminaries
In this section, we recall the concepts of metric mean dimensions in both topological and measure-theoretic situations, and state the precise definition of specification property in dynamical systems.
@@ 2.1 Three types of metric mean dimensions
Let $(X,d,f)$ be a dynamical system. For $n\in\mathbb{N}$, we define the Bowen metric $d_{n}$ on $X$ as
@@
$h_{top}^{UC}(f,Z,\epsilon)=\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(Z,\epsilon).$
Analogous to the definition of box dimension in fractal geometry, we can formulate the definition for upper capacity metric mean dimension of subsets using separated sets.
@@
When $Z=X$, we write $h_{top}(f,X,\epsilon)=h_{top}^{UC}(f,X,\epsilon)$, and $\overline{\rm mdim}_{M}(f,X,d)=\overline{\rm mdim}_{M}^{UC}(f,X,d)$, respectively. In this case, it is exactly reduced to Lindenstrauss and Weiss’s classical notion introduced in [LW00]. Such a formulation allows us to extend the metric mean dimension to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
Recall that the upper capacity topological entropy of $Z$ is limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 00, that is,
$h_{top}^{UC}(f,Z)=\lim_{\epsilon\to 0}h_{top}^{UC}(f,Z,\epsilon)=\sup_{\epsilon>0}h_{top}^{UC}(f,Z,\epsilon).$
Especially, for any system admitting the infinite topological entropy, topological entropy does not provide the additional information of the dynamics except we have known the system has pretty complicated topological complexity. Thus, (upper capacity) metric mean dimension can be viewed as a quantity to detect how fact the $\epsilon$-upper capacity topological entropy diverges to the infinite (upper capacity) topological entropy as $\epsilon$ tends to 00. More precisely, for sufficiently $\epsilon>0$, we may think of
@@
It is easy to see that any finite entropy system has zero metric mean dimension, and hence verifies that metric mean dimension is a good candidate to capture the dynamics of infinite entropy systems.
@@
The quantity $M(Z,s,N,\epsilon)$ is non-decreasing as $N$ increases. So we set
@@
Since the quantity $P(Z,s,N,\epsilon)$ is non-increasing as $N$ increases, and thus the following limit exists:
@@
One can replace $\limsup_{\epsilon\to 0}$ by $\liminf_{\epsilon\to 0}$ in above definitions to obtain the corresponding three types of lower metric mean dimensions on subsets, which are denoted by $\underline{\rm mdim}_{M}^{S}(f,Z,d)$ with $S\in\{UC,B,P\}$. If $\underline{\rm mdim}_{M}^{S}(f,Z,d)=\overline{\rm mdim}_{M}^{S}(f,Z,d)$, we call the common value ${\rm mdim}_{M}^{S}(f,Z,d)$ the (capacity/Bowen/packing) metric mean dimension of $Z$. We summarize some fundamental properties of metric mean dimensions derived in [YCZ22, Proposition 3.4] and [YCZ23b, Proposition 2.6].
@@
and $h_{top}^{UC}(f,Z,\epsilon)=\max_{1\leq i\leq N}h_{top}^{UC}(f,Z_{i},\epsilon)$ if $Z=\bigcup_{i=1}^{N}Z_{i}$.
@@
$(3)$ If $Z$ is a nonempty subset of $X$, then for every $10<\epsilon<$ 1, we have
$h_{top}^{B}(f,Z,3\epsilon)\leq h_{top}^{P}(f,Z,\epsilon)\leq h_{top}^{UC}(f,Z,\epsilon).$
@@
The metric mean dimension of a given subset is in general not easy to calculate its precise value. However, the Entropy Distribution Principle allows us to get a lower bound for $\epsilon$-Bowen and $\epsilon$-packing topological entropies of some fractal-like sets in multifractal analysis of dynamical systems. Consequently, the powerful tool gives the lower bounds of entropy and metric mean dimension of the fractal-like sets.
@@
[BR23, Lemma 13] Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Let $Z$ be a Borel set of $X$ and $s\geq 0$. If there exist a Borel probability measure $\mu\in M(X)$ and a constant $C>0$ satisfying:
@@
Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Let $Z$ be a Borel set of $X$ and $s\geq 0$. If there exist a Borel probability measure $\mu\in M(X)$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ of positive integer that converges to $\infty$ satisfying:
@@
holds for any Borel subset $E\subset Z$ with $\mu(E)>0$. For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of $E$. Then by $5r$-covering lemma stated in [Mat95, Theorem 2.1], there is a countable pairwise disjoint subcollection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, and in particular the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint, such that